OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 5

pooja195 (pooja195):

Hi welcome to openstudy :-) Use the following: \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

Im not new just a new account lol but thanks but I dont understand your response

pooja195 (pooja195):

Ok are you fimiliar with the point slope formula?

OpenStudy (anonymous):

Yes and no

pooja195 (pooja195):

What do you mean? What part don't you understand? :-/

OpenStudy (anonymous):

I do know about slope

pooja195 (pooja195):

Ok We have this formula called the point slope formula \[\huge~\rm~\bf~y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

ok

pooja195 (pooja195):

M is the slope what is the # of the slope given?

OpenStudy (anonymous):

8,3

OpenStudy (anonymous):

m=5

pooja195 (pooja195):

Good :) \[\huge~\rm~\bf~y-y_1=5(x-x_1)\] can you tell me what x=? y=?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Look at the coordinates

OpenStudy (anonymous):

8,3

pooja195 (pooja195):

Ok from that which number is x which is y

OpenStudy (anonymous):

idk

pooja195 (pooja195):

in a coordinate pair we have (x,y) (8,3) is our coordinate based off of this tell me what x and y are.

OpenStudy (anonymous):

idk tbh

pooja195 (pooja195):

x=8 because it is the first number y=3 because it is the 2nd number (x,y) (8,3) Now all you need to do is plug the numbers into the formula x1=8 y1=3

OpenStudy (anonymous):

Lol Im so confused

pooja195 (pooja195):

What is confusing?

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

Look Nvm

pooja195 (pooja195):



pooja195 (pooja195):

\[\huge~\rm~\bf~y-y_1=5(x-x_1)\] y1=3 x1=8 plug them in